OpenStudy (anonymous):

eg a11a22a33-a11a32a23-a12a21a33+a12a31a23+a13a21a32-a13a31a22 that...my teacher didnt quite get through to me with this one when do u use this in finding determinant?she explained it as something separate from cofactors so idk

OpenStudy (amistre64):



OpenStudy (amistre64):

i think that is what your refering to

OpenStudy (amistre64):

\begin{pmatrix} x&y&z\\ n&1&2\\ m&3&4 \end{pmatrix} lets try to use this instead, might be easser to read and keep track of

OpenStudy (amistre64):

\begin{pmatrix} x&&\\ &1&2\\ &3&4 \end{pmatrix} x(1*4-3*2) lets move the n & m over to help out the pattern \begin{pmatrix} &y&\\ &&2&n\\ &&4&m \end{pmatrix} -y(2m-4n) \begin{pmatrix} &&z\\ n&1&\\ m&3& \end{pmatrix} z(3n-1m) these are the parts of the overall determinate of the 3x3 matrix add them together to get the whole thing

OpenStudy (amistre64):

now, those arent to be regarded as numbers; only placeholders for clarity

OpenStudy (anonymous):

? this is my face right now o_O

OpenStudy (anonymous):

firstly it is not a cofactor method right? or ero? but u still multiply along main diagonal?

OpenStudy (amistre64):

\begin{pmatrix}\Large a_{11}&a_{12}&a_{13}\\ a_{21}&a_{22}&a_{23}\\ a_{31}&a_{32}&a_{33} \end{pmatrix} \[\Large D = a_{11}(a_{22}a_{33}-a_{32}a_{23})-...+...\] if your looking for the determinate of a 3x3 matrix; its alot to try to type out :)

OpenStudy (amistre64):

what is a cofactor method? its been awhile

OpenStudy (anonymous):

yh i get that part that u just typed was cofactor but the method in my question seems like cofactor

OpenStudy (amistre64):

the method in your example, is the same from what I can parse

OpenStudy (amistre64):

a11a22a33-a11a32a23-a12a21a33+a12a31a23+a13a21a32-a13a31a22 a11(a22a33-a32a23) -a12(a21a33-a31a23) +a13(a21a32-a31a22) oy thats hard on the eyes

OpenStudy (anonymous):

look at that slide 66 there about

OpenStudy (amistre64):

write, since the det is not 0, it is possible that there is an inverse to be found

OpenStudy (amistre64):

the next test is to see if they are linearly independant i believe

OpenStudy (amistre64):

or maybe not lol; the next slide has an adjoin that I am not familiar with yet